# MATHWELL: Generating Educational Math Word Problems at Scale

### v3

@@ C.1 Logistic Regression for Predicting MaC
Constant | 1.237 | 0.286 | 4.327 | 0.000** |
LLEMMA | -2.215 | 0.278 | -7.954 | 0.000** |
MAmmoTH | -0.855 | 0.241 | -3.548 | 0.000** |
Llama-2 | -0.492 | 0.245 | -2.006 | 0.045* |
Addition | 0.187 | 0.195 | 0.956 | 0.339 |
Subtraction | 0.439 | 0.207 | 2.124 | 0.034* |
Multiplication | 0.215 | 0.212 | 1.014 | 0.311 |
Division | -0.155 | 0.260 | -0.594 | 0.552 |
Fractions | -0.347 | 0.361 | -0.960 | 0.337 |
Decimals | -0.211 | 0.281 | -0.752 | 0.452 |
C.2 Accuracy by Question Type
As shown in Table 8, MATHWELL’s accuracy does not differ significantly or substantively by operation, as its accuracy remains above 90% for all operations, while the other models have a significant and/or substantive gap in their accuracy for the operation they perform best on relative to the operation they perform worst on. While MAmmoTH outperforms MATHWELL for addition, multiplication, and division, MATHWELL performs better in the other three operations and in overall accuracy.
Model | Add. | Sub. | Mult. | Div. | Frac. | Dec. |
LLEMMA | 76.2 | 72.3 | 63.4 | 56.0 | 50.0 | 63.2 |
MAmmoTH | 96.5 | 96.3 | 96.8 | 97.6 | 87.5 | 91.3 |
Llama-2 | 89.3 | 91.1 | 87.5 | 80.0 | 82.4 | 75.0 |
MATHWELL | 96.1 | 97.4 | 94.5 | 91.3 | 100.0 | 94.1 |
C.3 FKGL for All Versus MaC Questions
@@
C.4 Additional Automatic Evaluation Comparisons
@@
As shown in Table 9, PPL does not significantly vary for all models except for LLEMMA when comparing all to MaC generations. In the table, we also compare the BERTScore between all and MaC questions from each model and show they are similar. While LLEMMA has a longer average token length for all its generations than MATHWELL, MATHWELL’s MaC average token length is the longest of the models considered.
Model | PPL | M PPL | BF1 | M BF1 | A/M BF1 | G BF1 | AL | M AL | NDC |
LLEMMA | 3.82 (0.10) | 3.14 (0.10) | 84.3 | 85.3 | 84.6 | 84.1 | 56.4 (1.44) | 50.9 (2.89) | 8.41 (0.09) |
MAmmoTH | 2.76 (0.03) | 2.74 (0.04) | 86.0 | 86.4 | 86.1 | 84.7 | 45.9 (1.13) | 44.4 (1.15) | 8.25 (0.08) |
Llama-2 | 2.52 (0.03) | 2.52 (0.04) | 85.5 | 85.8 | 85.6 | 84.3 | 51.6 (0.99) | 49.8 (1.19) | 8.20 (0.07) |
MATHWELL | 2.44 (0.03) | 2.43 (0.03) | 85.5 | 85.7 | 85.6 | 84.2 | 54.7 (0.86) | 54.1 (0.97) | 8.23 (0.08) |
Appendix D Prompting Process
D.1 Standard Prompt
@@
Our standard prompt for interacting with the MATHWELL model is, "Write a grade school math word problem about {topic} and Python function with a commented out step-by-step solution to solve the word problem." In this prompt, topic is an optional argument, which we randomly select from a list of topics relevant to K-8 students. We begin every prompt with a random selection of 8-shot examples from SGSM Train.
D.2 Suggested Prompting Strategies
@@
We find MATHWELL is more likely to generate executable code when given a topic than when a topic is not specified. For example, when prompting our finetuned Llama-2 model before further training it on the SGSM Train data, we found the model generated executable code 63.1% of the time when given a topic, and only 32.3% of the time when a topic was not specified. As a result, for evaluating models in this paper, we provide them with a randomly selected topic, which also gives us the ability to assess their ability to effectively generate topic-specific word problems. Additionally, this evaluation strategy is aligned with how a teacher or student would use the model in practice, as they would want the generated questions to align with a particular topic. Qualitative evaluations of model generations also revealed that MATHWELL is more likely to generate executable code when the topic is more specific. For example, if their desired topic is superheroes, a user would have a higher likelihood of receiving a generation with executable code by prompting with a specific superhero (e.g., Superman) than leaving the topic general (e.g., superheroes).
Appendix E Example Generations
@@
E.1 LLEMMA
E.1.1 Solvability
@@ E.2.2 Accuracy
222MAmmoTH is prone to ending its generations with print(solution()) because its PoT training data always ends with a print statement Yue et al. (2023).
@@ E.4.3 Appropriateness
Appendix F Early MATHWELL Experimentation
In addition to training context-free question/answer pair generators, we also experiment with training context-free question generation models. Our theory is that if we could train a model to generate questions effectively, we could pass those questions to a math QA model to retrieve answers automatically. To test this theory, we finetune both Llama-2 and MAmmoTH as question generators using the same training data discussed in Section 3.1, except for excluding the solution for each question and modifying the standard prompt to ask the model to generate a question only. We then sample and evaluate 100 generations from each model. We find that MAmmoTH performs better than Llama-2 at this task, but neither model performs optimally. For example, only 19% of the MAmmoTH generations include the requested topic and 52.6% are solvable. Therefore, based on the results we report in Table 2, we conclude that it is more efficient to train a context-free question/answer pair generator than question generator.
Appendix G Annotation Process
G.1 Annotators
@@
G.2 Inter-Annotator Agreement
@@
G.3 Handling Annotator Disagreement
@@
G.4 Validating Final Evaluation Labels
For annotating the 250 samples from each model for our experiments reported in Section 4, we randomized questions from each model and had them blindly reviewed by one of our highly trained annotators with K-12 teaching experience. To evaluate the quality of these labels, we had 285 randomly reviewed by one additional annotator and 60 random reviewed by two additional annotators. The annotators agreed on solvability of the time, accuracy of the time, appropriateness of the time, all three labels of the time, and MaC of the time. Similar to above, the agreement rates for accuracy and solvability are higher than reported in recent studies that explore human alignment of LLM outputs, and the agreement rates for appropriateness and MaC are on par with these studies Bai et al. (2022); Ouyang et al. (2022); Stiennon et al. (2022); Ziegler et al. (2020).
Additional analysis reveals that most annotator disagreement (81.6%) was due to the primary annotator being more conservative than the additional annotators by labeling questions as not having the desired criteria when the additional annotators rated them as having the desired criteria. As a result, we chose to use the labels from the primary annotator when reporting final results to be conservative, though we also found the results do not vary when switching labels based on annotator disagreement. Annotators were least likely to disagree on labels for MATHWELL outputs and our primary annotator was not more likely to rate MATHWELL outputs as having the desired criteria than the additional annotators. Taken together, this evidence suggests our final labels are highly accurate.
G.5 Annotation Interface

### v4

@@ D.1 Logistic Regression for Predicting MaC
| Constant | 1.496 | 0.254 | 5.895 | 0.000** |
| GPT-4 Turbo | 0.064 | 0.262 | 0.243 | 0.808 |
| GPT-3.5 Turbo | -0.496 | 0.247 | -2.008 | 0.045* |
| LLEMMA | -2.279 | 0.276 | -8.269 | 0.000** |
| MAmmoTH | -0.867 | 0.238 | -3.638 | 0.000** |
| Llama-2 | -0.532 | 0.244 | -2.183 | 0.029* |
| Addition | 0.130 | 0.153 | 0.850 | 0.395 |
| Subtraction | 0.234 | 0.165 | 1.413 | 0.158 |
| Multiplication | -0.123 | 0.165 | -0.748 | 0.454 |
| Division | -0.124 | 0.193 | -0.645 | 0.519 |
| Fractions | -0.152 | 0.323 | -0.471 | 0.638 |
| Decimals | -0.345 | 0.208 | -1.658 | 0.097 |
D.2 Accuracy by Question Type
As shown in Table 7, MATHWELL’s accuracy does not differ significantly or substantively by operation, as its accuracy remains above 90% for all operations, while the other models except GPT-4 have a significant and/or substantive gap in their accuracy for the operation they perform best on relative to the operation they perform worst on. While MAmmoTH outperforms MATHWELL for addition, multiplication, and division, MATHWELL performs better in the other three operations and in overall accuracy. Further, MATHWELL outperforms GPT-4 in addition, subtraction, and decimals while maintaining similar performance in the other three operations.
| Model | Add. | Sub. | Mult. | Div. | Frac. | Dec. | |
| API | GPT-4 Turbo | 96.0 | 94.3 | 95.9 | 93.3 | 100.0 | 90.9 |
| GPT-3.5 Turbo | 91.0 | 95.2 | 86.6 | 87.5 | 71.4 | 90.2 | |
| Public | LLEMMA | 76.2 | 72.3 | 63.4 | 56.0 | 50.0 | 63.2 |
| MAmmoTH | 96.5 | 96.3 | 96.8 | 97.6 | 87.5 | 91.3 | |
| Llama-2 | 89.3 | 91.1 | 87.5 | 80.0 | 82.4 | 75.0 | |
| MATHWELL (Ours) | 96.1 | 97.4 | 94.5 | 91.3 | 100.0 | 94.1 |
D.3 FKGL for All Versus MaC Questions
@@
D.4 Additional Automatic Evaluation Comparisons
@@
As shown in Table 8, PPL does not significantly vary for all models except for LLEMMA when comparing all to MaC generations. In the table, we also compare the BERTScore between all and MaC questions from each model and show they are similar. While LLEMMA has a longer average token length for all its generations than MATHWELL, MATHWELL’s MaC average token length is the longest of the open-source models considered.
| Model | PPL | M PPL | BF1 | M BF1 | A/M BF1 | G BF1 | AL | M AL | NDC | |
| API | GPT-4 Turbo | 2.50 (0.03) | 2.49 (0.03) | 85.4 | 85.5 | 85.5 | 84.6 | 66.0 (2.12) | 66.8 (2.62) | 8.26 (0.07) |
| GPT-3.5 Turbo | 2.64 (0.03) | 2.70 (0.04) | 85.6 | 85.8 | 85.7 | 84.7 | 52.8 (1.00) | 49.9 (1.16) | 8.58 (0.07) | |
| Public | LLEMMA | 3.82 (0.10) | 3.14 (0.10) | 84.3 | 85.3 | 84.6 | 84.1 | 56.4 (1.44) | 50.9 (2.89) | 8.41 (0.09) |
| MAmmoTH | 2.76 (0.03) | 2.74 (0.04) | 86.0 | 86.4 | 86.1 | 84.7 | 45.9 (1.13) | 44.4 (1.15) | 8.25 (0.08) | |
| Llama-2 | 2.52 (0.03) | 2.52 (0.04) | 85.5 | 85.8 | 85.6 | 84.3 | 51.6 (0.99) | 49.8 (1.19) | 8.20 (0.07) | |
| MATHWELL (Ours) | 2.44 (0.03) | 2.43 (0.03) | 85.5 | 85.7 | 85.6 | 84.2 | 54.7 (0.86) | 54.1 (0.97) | 8.23 (0.08) |
Appendix E Finetuning Ablation
| Model | Solv. | Acc. | App. | MaC | EC | EC/MaC |
| MATHWELL (1st Stage) | 82.3 (0.67) | 85.8 (0.66) | 66.0 (0.83) | 58.9 (0.87) | 63.1 (0.73) | 37.2 (0.55) |
| MATHWELL (2nd Stage) | 89.2* (1.97) | 96.9* (1.17) | 86.5* (2.29) | 74.8* (2.75) | 66.4* (1.00) | 49.6* (1.83) |
As shown in Table 9, MATHWELL shows a statistically and substantively meaningful improvement in all metrics after the second stage of finetuning on high-quality synthetic data labeled by domain experts. These results lead us to conclude that the second stage of finetuning is critical for improving the model’s ability to generate questions that are educationally appropriate.
Appendix F Question Classifier Performance
| Solvability | Accuracy | Appropriateness | ||||||||||||||||
| Annotated Test | 78.9 | 61.9 | 86.7 | 88.0 | 87.4 | 0.701 | 81.6 | 53.8 | 87.2 | 92.2 | 89.6 | 0.649 | 82.7 | 72.0 | 89.8 | 88.9 | 89.4 | 0.793 |
| GPT-4 Turbo | 83.6 | 62.3 | 96.2 | 86.1 | 90.9 | 0.560 | 95.8 | 54.8 | 96.2 | 99.6 | 97.8 | 0.637 | 77.2 | 59.0 | 87.2 | 85.5 | 86.4 | 0.579 |
| GPT-3.5 Turbo | 81.9 | 52.3 | 88.5 | 91.3 | 89.9 | 0.519 | 88.6 | 55.2 | 90.5 | 97.4 | 93.9 | 0.541 | 78.1 | 64.3 | 82.1 | 91.0 | 86.3 | 0.696 |
| LLEMMA | 48.8 | 49.8 | 48.7 | 90.2 | 63.2 | 0.542 | 59.0 | 47.1 | 62.5 | 89.7 | 73.7 | 0.492 | 45.1 | 50.1 | 41.8 | 80.4 | 55.0 | 0.489 |
| MAmmoTH | 84.4 | 52.5 | 87.4 | 95.9 | 91.4 | 0.700 | 93.5 | 57.9 | 95.7 | 97.6 | 96.6 | 0.669 | 66.8 | 54.2 | 69.8 | 89.8 | 78.6 | 0.574 |
| Llama-2 | 82.4 | 53.1 | 84.9 | 96.2 | 90.2 | 0.619 | 85.2 | 51.6 | 89.8 | 94.1 | 91.9 | 0.677 | 73.8 | 50.4 | 81.1 | 88.2 | 84.5 | 0.628 |
| MATHWELL | 85.2 | 54.3 | 90.1 | 93.7 | 91.9 | 0.599 | 94.6 | 62.7 | 97.7 | 96.8 | 97.2 | 0.790 | 77.6 | 63.1 | 90.4 | 82.9 | 86.5 | 0.669 |
As shown in Table 10, despite weighting the loss, each classifier has a lower balanced accuracy than its overall accuracy. The classifiers perform similarly on outputs from all models except LLEMMA, for which performance is significantly lower. As would be expected, performance for all models tends to be lower than on the test dataset. Excluding LLEMMA and the appropriateness classifier’s precision on MAmmoTH outputs, each classifier has fairly high precision, recall and F1, suggesting they are effectively able to label solvable, accurate, and appropriate questions. The classifiers tend to have low ROC AUC across sources, suggesting they do not perform well at different prediction cutoffs. As a whole, Table 10 provides evidence that text classifiers can learn some features that are important for labelling MATHWELL outputs, but future research should use more balanced datasets to improve their ability to label unsolvable, inaccurate, and inappropriate questions.
Appendix G Topics for Data Generation
We collected topics and keywords in collaboration with our volunteer annotators with K-12 teaching experience. The topics and keywords span a wide array of student interests including sports, music, video games, movies/TV, animals, vehicles, and food. On top of being manually collected from experts, we believe the resultant list is largely inarguable.
We used a smaller list of topics (43 total) when comparing models for the experiments reported in Section 4 to make sure the generations could be compared fairly and followed a similar contextual distribution. When generating SGSM, we used a much larger list of topics (185) and included generations that did not have a specified topic (21.4% of samples) to promote topical diversity. Based on the automatic evaluations reported in Section 4.2 and initial qualitative evaluation of SGSM, we believe the dataset’s characteristics are equivalent to those of the MATHWELL data evaluated by human experts in Section 4 despite being based on a larger list of topics. The full lists of topics written in Python list format are below.
Topics Used to Compare Models
: [’Superman’, ’Batman’, ’Wonder Woman’, ’Barbie’, ’Power Rangers’, ’basketball’, ’soccer’, ’football’, ’volleyball’, ’field hockey’, ’Fortnite’, ’Spiderman’, ’Iron Man’, ’Captain America’, ’Captain Marvel’, ’Thor, the God of Thunder’, ’Ninja Turtles’, ’Black Panther’, ’Taylor Swift’, ’swimming’, ’Pokémon’, ’Super Mario’, ’Naruto’, ’unicorns’, ’Hello Kitty’, ’Minecraft’, ’lacrosse’, ’cheer leading’, ’LeBron James’, ’Steph Curry’, ’Patrick Mahomes’, ’Serena Williams’, ’dogs’, ’cats’, ’dinosaurs’, ’Harry Potter’, ’cars’, ’planes’, ’trains’, ’pizza’, ’cookies’, ’ice cream’, ’candy’]
Topics Used to Generate SGSM
: [’Famous Artists’, ’Detectives and Mystery’, ’Cartoons and Animated Shows’, ’Famous Landmarks’, ’Pets (cats, dogs, hamsters)’, ’Beauty and the Beast’, ’Underwater Creatures’, ’Music Instruments’, ’cheer leading’, ’Pottery and Ceramics’, ’Famous Scientists’, ’Fashion and Clothing’, ’lacrosse’, ’Zoo Animals’, ’the Circus and Performing Arts’, ’The Lion King’, ’Harry Potter’, ’Health and Nutrition’, ’DIY Science Projects’, ’Geology’, ’Cooking and Baking’, ’cars’, ’SpongeBob SquarePants’, ’Alice in Wonderland’, ’Space and Astronomy’, ’Magic Tricks’, ’Scooby-Doo’, ’Ratatouille’, ’Paw Patrol’, ’Animals and Wildlife’, ’Unexplained Mysteries’, ’Puzzles’, ’Insects and Bugs’, ’Famous Inventors’, ’Fairy Gardens’, ’cookies’, ’swimming’, ’basketball’, ’Minions’, ’Ancient Civilizations’, ’Weather and Meteorology’, ’Virtual Reality’, ’Fairies’, ’Environmental Conservation’, ’Mummies’, ’Art and Drawing’, ’The Avengers’, ’Dora the Explorer’, ’dogs’, ’Historical Figures’, ’ice cream’, ’Space Exploration’, ’unicorns’, ’Mythical Creatures’, ’Aliens’, ’Pocahontas’, ’Peppa Pig’, ’The Simpsons’, ’soccer’, ’Ancient Egypt’, ’Taylor Swift’, ’Up’, ’Pikachu’, ’Winnie the Pooh’, ’Construction and Building’, ’Firefighters’, ’Fairy Tales and Folklore’, ’Friendship and Relationships’, ’field hockey’, ’Sesame Street’, ’The Incredibles’, ’Looney Tunes’, ’Batman’, ’Riddles’, ’Natural Disasters’, ’Patrick Mahomes’, ’Ocean Life’, ’Wildlife Conservation’, ’Cars and Racing’, ’Captain Marvel’, ’Sonic the Hedgehog’, ’Marine Biology’, ’Farm Life’, ’Myths and Legends’, ’Robots and Technology’, ’Countries and Cultures’, ’Doctors and Nurses’, ’Hiking and Nature’, ’Hello Kitty’, ’Ninja Turtles’, ’Pokémon’, ’Whales and Dolphins’, ’LeBron James’, ’Black Panther’, ’Science Experiments’, ’Science Fiction’, ’Fortnite’, ’planes’, ’The Justice League’, ’Hiking and Camping’, ’Robots’, ’Barbie’, ’Cultural Festivals’, ’Inventions’, ’Toy Story’, ’Time Travel’, ’Photography’, ’Ballet and Dance’, ’Wizards’, ’volleyball’, ’Disney Princesses’, ’Power Rangers’, ’Astronauts and Space Exploration’, ’dinosaurs’, ’Minecraft’, ’Camping’, ’Famous Explorers’, ’Computer Programming’, ’Peter Pan’, ’Spiderman’, ’candy’, ’Moana’, ’Frozen (Elsa and Anna)’, ’Reptiles and Amphibians’, ’Space and Astronauts’, ’Dance and Ballet’, ’Transformers’, ’Naruto’, ’Plants and Gardening’, ’Cars and Vehicles’, ’Famous Composers’, ’Jurassic Park’, ’Indoor Games’, ’Super Mario’, ’Steph Curry’, ’Frozen’, ’Star Wars’, ’Space Aliens and UFOs’, ’Iron Man’, ’Planets and the Solar System’, ’Finding Nemo’, ’Forests and Trees’, ’Pirates’, ’National Parks’, ’Police Officers’, ’Volcanoes’, ’Superman’, ’Travel and Adventure’, ’Serena Williams’, ’Cinderella’, ’DIY Crafts’, ’Exploring the Ocean’, ’Outer Space’, ’Magic and Wizards’, ’cats’, ’Mythical Creatures (dragons, unicorns)’, ’Thor, the God of Thunder’, ’Farm Animals’, ’The Solar System’, ’Captain America’, ’Underwater Life’, ’Recycling’, ’Tom and Jerry’, ’Birds and Bird Watching’, ’The Jungle Book’, ’Books and Reading’, ’Aladdin’, ’pizza’, ’Board Games’, ’Construction Vehicles’, ’My Little Pony’, ’Cars’, ’Native American Culture’, ’Mythology’, ’Pirates and Treasure’, ’Shrek’, ’Music and Instruments’, ’Mystery Stories’, ’The Little Mermaid’, ’Wonder Woman’, ’Games and Puzzles’, ’trains’, ’Circus Animals’, ’football’, ’Thomas the Tank Engine’]
Appendix H Prompting Process
H.1 Standard Prompt
@@
Our standard prompt for interacting with the MATHWELL model is, "Write a grade school math word problem about {topic} and Python function with a commented out step-by-step solution to solve the word problem." In this prompt, topic is an optional argument, which we randomly select from the list of topics discussed in Appendix G. We begin every prompt with a random selection of 8-shot examples from SGSM.
H.2 Suggested Prompting Strategies
@@
We find MATHWELL is more likely to generate executable code when given a topic than when a topic is not specified. For example, when prompting our finetuned Llama-2 model before further training it on the SGSM data, we found the model generated executable code 63.1% of the time when given a topic, and only 32.3% of the time when a topic was not specified. As a result, for evaluating models in this paper, we provide them with a randomly selected topic, which also gives us the ability to assess their ability to effectively generate topic-specific word problems. Additionally, this evaluation strategy is aligned with how a teacher or student would use the model in practice, as they would want the generated questions to align with a particular topic. Qualitative evaluations of model generations also revealed that MATHWELL is more likely to generate executable code when the topic is more specific. For example, if their desired topic is superheroes, a user would have a higher likelihood of receiving a generation with executable code by prompting with a specific superhero (e.g., Superman) than leaving the topic general (e.g., superheroes).
Appendix I GPT-4 Coding Performance
GPT-4 tends to output additional information beyond what is requested in the prompt and shown in few-shot examples, such as example usage or comments about the Python function solution (see Appendix J.1.2 for an example). This results in the model having a lower percentage of executable code than GPT-3.5 when its output is parsed by the same script we used to parse the output from all other models we evaluated. We used the same parsing script across models to evaluate them all the same way, rather than creating a customized script for GPT-4.
Appendix J Example Generations
@@ Appendix I GPT-4 Coding Performance
J.1 GPT-4 Turbo
J.1.1 Solvability
Positive Example
In a volleyball match, Team A scored twice as many points as Team B. If Team B scored 15 points, how many points in total were scored in the match?
Negative Example
In a Fortnite Battle Royale match, there are 100 players. If one player can eliminate 4 opponents, how many players in the game can achieve such an elimination count, assuming no other players are eliminated?
Negative Example Reasoning
It is not entirely clear what this question is asking. If it defined how many players should be left in the game after eliminations, it would be possible to determine how many players can be eliminated but, without defining that, it is not clear what the student needs to do to solve this problem.
J.1.2 Accuracy
Positive Example
Jenny has 15 cats. She buys 2 new cat towers, each can hold 4 cats. If she places as many cats as possible on the cat towers, how many cats are not on the towers?
Solution:
def solution():
# Jenny starts with 15 cats
total_cats = 15
# Each new cat tower holds 4 cats
tower_capacity = 4
# She buys 2 new cat towers
total_towers = 2
# The total number of cats the towers can hold is
total_capacity = tower_capacity * total_towers
# The number of cats not on the towers is
cats_not_on_towers = total_cats - total_capacity
return cats_not_on_towers
# Uncomment the line below to test the function with the given values
# print(solution())444As discussed in Appendix I, GPT-4 tends to output additional information beyond what is requested in the prompt and shown in few-shot examples, such as example usage or comments about the Python function solution, as is included here.
Negative Example
Mrs. Johnson has 12 adult cats and each adult cat has 3 kittens. After several months, those kittens grow up and each new adult cat also has 2 kittens. How many kittens are there in total now?
Solution:
def solution():
# Mrs. Johnson has 12 adult cats originally
adult_cats_initial = 12
# Each adult cat has 3 kittens
kittens_per_adult_initial = 3
# Calculate the initial number of kittens
initial_kittens = adult_cats_initial *
kittens_per_adult_initial
# The initial kittens grow up to become adults
new_adult_cats = initial_kittens
# Each new adult cat has 2 kittens
kittens_per_new_adult = 2
# Calculate the new kittens born from the
# grown-up kittens
new_kittens = new_adult_cats *
kittens_per_new_adult
# The total number of kittens now is the sum
# of the initial kittens and new kittens
total_kittens = initial_kittens + new_kittens
return total_kittens
# Call the function and print the result
print(f"There are {solution()} kittens in total now.")
Negative Example Reasoning
The solution adds the original number of kittens to the number of new kittens, which is incorrect because the original kittens are now considered adults based on the text of the question.
J.1.3 Appropriateness
Positive Example
During a week of intense battles, the Power Rangers defeated 15 evil monsters on Monday, twice as many on Tuesday as on Monday, and 10 fewer on Wednesday than on Tuesday. How many monsters did they defeat in total from Monday to Wednesday?
Negative Example
Serena Williams has won a certain number of tennis matches this season. If she wins 8 more matches, she will double the number she has won so far. How many matches has Serena won this season?
Negative Example Reasoning
This question does not require any mathematical operation to solve and is therefore not appropriate. The question tells you Serena’s number of wins will double if she wins 8 more matches, which means she had to have won 8 matches so far.
J.2 GPT-3.5 Turbo
J.2.1 Solvability
Positive Example
During a basketball game, the Lakers made a total of 84 points. LeBron James scored 35 points, and Anthony Davis scored 20 points. How many points were scored by the rest of the Lakers team?
Negative Example
The Ninja Turtles went to the pizza parlor and ordered 8 pizzas. Each pizza had 8 slices. If they each ate 2 slices, how many slices of pizza were left over?
Negative Example Reasoning
This question does not define the number of Ninja Turtles. While many students might know there are typically 4 main Ninja Turtles, not every student would know this and all math questions should still define the key variables you need to answer them.
J.2.2 Accuracy
Positive Example
Black Panther has 1500 vibranium shards. He needs to divide them equally amongst his 5 closest allies. How many vibranium shards will each ally receive?
Solution:
def solution():
#Black Panther has 1500 vibranium shards
vibranium_total = 1500
#He needs to divide them equally amongst his 5
#allies
allies = 5
#The answer is
result = vibranium_total / allies
return result
Negative Example
A dog shelter has 40 dogs. If each dog needs 2 cups of food per day, and a bag of food contains 40 cups, how many bags of food will the shelter need for a 30-day period?
Solution:
def solution():
#The shelter has 40 dogs
dogs = 40
#Each dog needs 2 cups of food per day
food_per_dog = 2
#A bag of food contains 40 cups
food_per_bag = 40
#For 30 days, the shelter will need
total_food_needed = dogs * food_per_dog * 30
#The answer is
result = total_food_needed // food_per_bag + 1
return result
Negative Example Reasoning
The solution adds an additional bag of food to the total assuming that the answer is a decimal and, therefore, that the shelter would need to buy another full bag since a partial bag is not possible. However, since the answer is whole number, this additional bag leads to outputting an incorrect result.
J.2.3 Appropriateness
Positive Example
A cheerleading team has 20 members. They want to evenly divide into 4 squads for a competition. How many members will be on each squad?
Negative Example
In a Pokémon battle, Pikachu has a 60% chance of winning each round. If Pikachu and Ash battle for 5 rounds, what is the probability that Pikachu wins all 5 rounds?
Negative Example Reasoning
This question is too hard for a middle school student. It assumes a student knows about the probability of independent events, which is typically not covered until high school or an introductory statistics college course.
J.3 LLEMMA
J.3.1 Solvability
@@
555MAmmoTH is prone to ending its generations with print(solution()) because its PoT training data always ends with a print statement Yue et al. (2023).
@@
Appendix K Early MATHWELL Experimentation
In addition to training context-free question/answer pair generators, we also experiment with training context-free question generation models. Our theory is that if we could train a model to generate questions effectively, we could pass those questions to a math QA model to retrieve answers automatically. To test this theory, we finetune both Llama-2 and MAmmoTH as question generators using the same training data discussed in Section A.1, except for excluding the solution for each question and modifying the standard prompt to ask the model to generate a question only. We then sample and evaluate 100 generations from each model. We find that MAmmoTH performs better than Llama-2 at this task, but neither model performs optimally. For example, only 19% of the MAmmoTH generations include the requested topic and 52.6% are solvable. Therefore, based on the results we report in Table 3, we conclude that it is more efficient to train a context-free question/answer pair generator than question generator.
Appendix L Annotation Process
L.1 Annotators
@@
L.2 Inter-Annotator Agreement
@@
L.3 Handling Annotator Disagreement
@@
L.4 Validating Final Evaluation Labels
For annotating the 250 samples from each model for our experiments reported in Section 4, we randomized questions from each model and had them blindly reviewed by one of our highly trained annotators with K-12 teaching experience. To evaluate the quality of these labels, we had 357 randomly reviewed by one additional annotator and 60 random reviewed by two additional annotators. The annotators agreed on solvability of the time, accuracy of the time, appropriateness of the time, all three labels of the time, and MaC of the time. Similar to above, the agreement rates for accuracy and solvability are higher than reported in recent studies that explore human alignment of LLM outputs, and the agreement rates for appropriateness and MaC are on par with these studies Bai et al. (2022); Ouyang et al. (2022); Stiennon et al. (2022); Ziegler et al. (2020).
Additional analysis reveals that most annotator disagreement (80.1%) was due to the primary annotator being more conservative than the additional annotators by labeling questions as not having the desired criteria when the additional annotators rated them as having the desired criteria. As a result, we chose to use the labels from the primary annotator when reporting final results to be conservative, though we also found the results do not vary when switching labels based on annotator disagreement. Annotators were least likely to disagree on labels for MATHWELL and GPT-4 Turbo outputs and our primary annotator was not more likely to rate MATHWELL outputs as having the desired criteria than the additional annotators. Taken together, this evidence suggests our final labels are highly accurate.
L.5 Annotation Interface